A capacitor charging circuit consists of a battery, an uncharged 20 μF capacitor, and a 5.0 kΩ resistor. At t = 0 s the switch is closed; 0.15 s later, the current is 0.38 mA . Part A What is the battery's emf in volts? Express your answer to two significant figures and include the appropriate units.
Answer #1

R= 5 k ohm

R= 5000 ohm

C=20uF= 20x10^-6

Time constant = Rc=5000(20x10^-6)=0.1 s

i= 0.38 mA= 0.38x10^-3 A

t= 0.15 s

i= (V/R) (e-t/RC)

0.38x10^-3 =(V/5x10^3)(e-0.15/0.1)

0.38x10^-3 =( V/5x10^3) 0.223

V/5x10^3 = 0.38x10^-3 /0.223

V=(1.7x10^-3 )(5x10^3)

V=8.5 volts.